How long will it take a 50 watts heater to melt 100 g of ice at 0 C
1
Expert's answer
2020-03-06T10:28:10-0500

Solution. Using the formula that relates heat energy to mass and heat of fusion:

"q=m \Delta H_f"

where q is heat energy; m=100g is mass; ∆Hf= 334 J/g is heat of fusion for ice. On the other hand, the heat generated by the heater is equal to


"q=Pt"

where P=50 watts is power; t is time. Since the energies are equal, we obtain


"m \Delta H_f=Pt"

Therefore


"t=\frac {m\Delta H_f}{P}=\frac{100\times 334}{50}=668s"

Answer. 668 seconds